# Data Readiness for AI: A 360-Degree Survey

### v1

@@ 4. DRAI Metrics
In this section, we describe the existing metrics found in the literature that are used to measure the preparedness of data for AI. We will explore both structured and unstructured data, with a primary focus on structured data. Nevertheless, we briefly touch upon important unstructured data metrics, including textual, multimedia, image, speed, and video-related data. In Table 3, we provide a summary of the dimensions and metrics defined in this study.
Table 3. Dimensions and Metrics for (Structured and Unstructured) Data Readiness for AI
@@
Santos et al. (Santos et al., 2020), Gupta et al. (Gupta et al., 2021)
@@
Templin (Templin, 1957), McCarthy et al. (mccarthy2007theoretical), McCarthy et al.(McCarthy and Jarvis, 2010)
@@
Rousseeuw et al. (Rousseeuw and Hubert, 2018), Degirmenci et al. (Degirmenci and Karal, 2021),
Gupta et al.’s (Gupta et al., 2021)
@@
Mislabels Gupta et al’s (Gupta et al., 2021), Cohen’s Kappa (Cohen, 1960) Readability Score
@@
Monge et al. (Monge and Elkan, 1996), Russell et al. (Russell, 1922),
Gupta et al. (Gupta et al., 2021)
@@
Hall et al. (Hall and Smith, 1999), Gupta et al. (Gupta et al., 2021)
@@
Gupta et al. (Gupta et al., 2021)
@@
Vatsalan et al. (Vatsalan et al., 2022), Duddu et al. (Duddu et al., 2022), Carlini et al. (Carlini et al., 2022),Song et al. (Song and Mittal, 2021), Bezzi (Bezzi, 2007), Longpr ’e et al. (Longpré et al., 2017), Sevgi et al. (Arca and Hewett, 2020)
@@
FAIRness Score Wilkinson et al. (Wilkinson et al., 2018), Clarke et al. (Clarke et al., 2019)
@@ 4.1. Structured data
This section aims to comprehensively define the various dimensions of readiness of structured data for AI and support the definitions with relevant examples. Additionally, we will emphasize the significance of existing metrics and scoring mechanisms presented in the literature that enable the measurement and evaluation of structured data readiness for AI across each dimension. It is important to note that specific dimensions discussed may be applicable primarily to supervised learning (e.g., mislabeled data), while others may be more relevant to unsupervised learning (e.g., duplicity). By exploring these dimensions and metrics within their respective learning paradigms, this paper contributes to a deeper understanding of the crucial factors in preparing data for successful AI training.
@@ 4.1.1. Completeness
Example: In a dataset containing information about customer demographics, if the “income” attribute is complete for all individuals, there are no missing values for “income” across the entire dataset. This completeness ensures that a dataset is reliable and suitable for analysis, as there is no loss of information due to missing data in the “income” attribute.
Metrics in Literature: Blake et al. (Blake and Mangiameli, 2011) propose the “completeness” metric to measure the presence of missing values in a dataset. This metric quantifies the proportion of data records with null (missing) values compared to the total number of data records. By conducting a thorough examination of the completeness metric, researchers demonstrate how the presence of missing data can significantly impact the results of classification tasks. For example, Jäger et al. (Jäger et al., 2021) demonstrates that handling missing values significantly enhances predictive model performance, resulting in substantial improvements of up to 20% for classification tasks and around 15% for regression tasks, emphasizing the importance of addressing missing data to optimize downstream machine learning outcomes.
In the context of addressing missing data, Santos et al. (Santos et al., 2020) emphasize the importance of data imputation, which involves filling in or estimating missing values to maintain data integrity. Particularly, the authors highlight the efficacy of the k-nearest neighbors (KNN) imputation technique in this process. They discuss the significance of choosing appropriate distance metrics, such as HVDM and HEOM, which effectively handle both nominal and continuous data while preserving data distribution during imputation. In line with this, Bors et al. (Bors et al., 2018) propose a different approach to quantify missing values in a dataset, using indicators to distinguish missing from non-missing values. Their method offers a practical tool for data preparation, enabling easy identification of missing data in AI applications. Synthesizing these research findings, we can extract valuable insights into effective data imputation and missing data handling, which collectively contribute to enhancing data quality and bolstering the trustworthiness of AI analysis and modeling.
Gupta et al. (Gupta et al., 2021) further contribute to the field with their data completeness metric in the data quality toolkit. This metric is designed to detect and handle missing values in datasets, which can significantly impact ML models and lead to inaccurate inferences. Valuable insights are provided by the quality component of the metric through its ability to identify the presence and location of missing values in the data. When the score is 1, it indicates that there are no missing values in the dataset. However, when the score is not perfect, the metric offers a remediation mechanism that leverages constraints and an association-based approach. This approach takes into consideration the diverse data types and the relationships between attributes, enabling the imputation of missing entries and enhancing the overall quality and reliability of the dataset for ML modeling purposes. Thus, Gupta et al.’s work complements Santos et al.’s findings by providing a practical and effective method for addressing missing data and improving the quality of datasets.
Gupta et al. (Gupta et al., 2021) contribute a data completeness metric in their toolkit, addressing missing values in datasets. The metric, part of the data quality toolkit, detects and handles missing values, offering insights into their presence and location. A perfect score indicates no missing values, while an imperfect score prompts a remediation mechanism. Leveraging constraints and an association-based approach, it considers data types and attribute relationships for effective imputation.
@@
Summary: Data completeness is crucial for accurate analysis and modeling. Metrics proposed by various researchers quantify the impact of missing values and suggest remedies such as KNN imputation with suitable evaluation metrics. Additionally, using indicators for distinction, and incorporating completeness metrics that consider data types and relationships further improve the handling of missing data.
@@ 4.1.2. Outliers
An outlier in a dataset refers to a data point or instance that significantly deviates from the typical or expected values within the dataset. It is an observation that lies far away from most of the data points and does not conform to the general patterns or trends present in the dataset.
@@
Example: Consider a dataset of housing prices based on factors such as location, size, and number of bedrooms. In this dataset, an outlier could be a property with an extremely high price that does not align with the average price range of similar properties. This outlier might be an exceptional case, such as a luxury mansion in an otherwise average neighborhood, or it could be a data entry error.
@@
Metrics in Literature: In their research, Bors et al. (Bors et al., 2018) discuss the concept of plausibility as a metric to identify outliers in datasets for AI applications, which can disrupt statistical analyses and modeling. To quantify the number of outliers, data analysts employ two main approaches: non-robust statistics and robust statistics. On the one hand, non-robust statistics involve using the mean and standard deviation to identify entries that deviate significantly from the mean. On the other hand, robust statistics use the median and the robust interquartile range estimator, which are more resistant to outliers. Li et al. (Li et al., 2021) further explore the standard deviation and interquartile range-based outlier detection methods in their study.
@@
Breunig et al. (Breunig et al., 2000) introduce the Local Outlier Factor (LOF) as a metric for identifying outliers in a dataset. LOF quantifies the level of being an outlier for each data instance by considering the density of the dataset’s distribution. Outliers are expected to have lower local densities compared to their surrounding instances. The LOF algorithm computes the LOF value for a specific instance by comparing the density of that instance’s neighborhood with its neighboring instances, where the neighborhood is defined by a user-defined parameter (e.g., number of nearest neighbors). Higher LOF values indicate a higher probability of an instance being an outlier, implying a notably lower density in its local neighborhood than neighboring data points. The LOF approach is flexible and adaptable, incorporating distribution density information to effectively identify outliers not detectable by other methods based solely on statistical measures or global properties of the data. Building on this foundation, Pokraja et al. (Pokrajac et al., 2007) introduced the ILOF (Incremental Local Outlier Factor) method, which uses the LOF (Local Outlier Factor) metric to determine if a new data point is an outlier. By analyzing the computed LOF value, the ILOF method assigns a score to the incoming sample, indicating whether it is classified as an outlier or not. This approach allows for real-time outlier detection and updates the scores of existing points to reflect the impact of the new data point.
@@
Degirmenci et al. (Degirmenci and Karal, 2021) introduce the RiLOF method, which addresses limitations in existing statistical outlier detection techniques by introducing the MoNNAD (Median of Nearest Neighborhood Absolute Deviation) metric. This metric is calculated as the median of the absolute variances among the LOF values of the $k$-nearest neighboring data points and the LOF value of a given sample. This score indicates how much the sample deviates from its local neighborhood. In the RiLOF method, the MoNNAD score is used to label and score query samples. Samples are categorized as outliers when their MoNNAD scores are equal to or greater than a specified threshold value. Unlike traditional techniques, the RiLOF method assigns more importance to the query sample, resulting in clearer differentiation between inliers and outliers. The study demonstrates that the MoNNAD metric, incorporated in the RiLOF method, successfully detects outliers, including outlier clusters, that other techniques fail to recognize.
@@
The GESD (Generalized Extreme Studentized Deviation) technique, as introduced by Rosner et al. (Rosner, 1983), and the MAD (Median Absolute Deviation) technique, proposed by Leys et al. (Leys et al., 2013), are both outlier detection methods that share similarities in their underlying principles. Both techniques aim to identify outliers within datasets by using statistical measures to assess the deviation of data points from central tendencies. GESD identifies outliers by evaluating the maximum absolute difference between each sample and the dataset’s mean, while normalizing it by the standard deviation. MAD identifies outliers by considering the median of absolute differences between data records and the dataset’s median, incorporating a constant associated with the assumption of data normality. Additionally, the Z-score method aligns with this principle by normalizing sample values using the mean and standard deviation or MAD. The robust Z-score version, introduced by Rousseeuw et al. (Rousseeuw and Hubert, 2018), substitutes the median and MAD for more resilient measures, demonstrating the shared concept of using statistical measures to detect outliers in datasets.
Gupta et al. (Gupta et al., 2021) contribute to outlier detection by introducing a new metric in their toolkit, which is designed to pinpoint and eliminate outliers within the dataset. The metric uses an approach for detecting outliers and provides an aggregate score between 0 and 1 to quantify the presence of outliers. A score of 1 indicates that there are no outliers. Additionally, the metric provides the exact indices of the outlier samples, facilitating easy removal and complementary remediation. By incorporating outlier detection, Gupta et al. offer a comprehensive solution for flagging and handling outliers, enhancing the quality and reliability of the dataset for improved model training.
@@
Summary: Outliers are data points that deviate significantly from most of the dataset and do not follow the general patterns. Different metrics and techniques have been proposed to identify and measure outliers in datasets. These include measures based on column heterogeneity, statistical measures like median, standard deviation, mean, and interquartile range, and techniques such as Local Outlier Factor (LOF), Generalized Extreme Studentized Deviation (GESD), Z-score.
@@ 4.1.3. Mislabeled Data
Mislabeled data in the context of preparing a dataset for AI refers to instances or data points with incorrect or inaccurate labels. It represents a form of labeling error or inconsistency within the dataset, where the assigned labels do not align with the true or expected labels.
@@
Metrics in Literature: Gupta et al.’s (Gupta et al., 2021) data quality toolkit introduces the label purity metric that evaluates the effect of inducing random noise on the performance of a classifier trained on the dataset. In the example provided in the study, $10\%$ random noise is introduced to $41$ datasets from UCI (Dua et al. (Dua and Graff, 2017)) and Kaggle ((Kag, nd)) repositories, and the performance of an AutoAI classifier (Liu et al. (Liu et al., 2020)) is measured using 3-fold cross-validation. The results show a drop in classifier performance after inducing noise, with varying degrees of decrease observed across the datasets. The label purity metric assigns a quality score to indicate the presence of noise in the data, where a score of $1$ represents an absence of noise. It also highlights the indices of noisy samples and suggests correct labels for them.
In evaluating the accuracy of labels assigned by multiple annotators, Cohen’s Kappa (Cohen, 1960) is widely employed for assessing inter-rater reliability, especially with categorical or binary labels. Cohen’s Kappa calculates the agreement beyond chance, ranging from -1 to 1, where values near 1 indicate substantial agreement. Lavitas et al. (Lavitas et al., 2021) contribute a credibility metric, assessing the likelihood of correct annotations based on multiple reviewers’ agreement. The metric ranges from $(N/2+1)/N$ to 1, reflecting high credibility with close agreement and lower credibility with less consensus among reviewers, offering insights into the reliability of the annotation process involving multiple reviewers.
@@ 4.1.4. Duplicity
Refers to the presence of duplicate or redundant instances or records within a dataset. It occurs when the same or very similar data entries appear multiple times, potentially skewing the analysis or modeling process.
Example: Consider a dataset of customer transactions where each entry represents a purchase made by a customer. Suppose there are duplicate records in the dataset. In that case, a particular transaction is represented more than once, leading to an over-representation of that specific transaction in the analysis. This duplication can distort the patterns, statistics, or relationships derived from the data, potentially affecting the accuracy and effectiveness of AI models trained on such a dataset.
Metrics in Literature: Bors et al. (Bors et al., 2018) propose a mechanism to identify duplicate entries in a dataset using a scoring system based on uniqueness. The user selects one or more columns in the dataset that are expected to have unique combinations of values. The system assigns a score of $1$ (true) to values with a unique combination in the selected columns and a 00 (false) score to values found multiple times. By incorporating this score, the system effectively flags duplicate entries in the dataset. This scoring system provides a way to distinguish between unique and duplicate values, contributing to data quality in AI applications where the accurate identification of duplicates is essential.
@@
Elmagarmid et al. (Elmagarmid et al., 2007) conducted a comprehensive survey exploring various similarity metrics for duplicate detection, addressing challenges in handling typographical variations in string data. One of the highlighted character-based similarity metrics in the survey is the Levenshtein distance metric (Levenshtein, 1965). This metric measures the number of operations needed to transform one string into another through edit operations (insertion, deletion, and character replacement). Waterman et al. (Waterman et al., 1976) introduced the Affine Gap Distance metric as a solution to the limitations of the standard edit distance metric in matching shortened or truncated strings. It introduces two edit operations: open and extend the gap. An open gap marks the beginning of a gap in sequence alignment, representing missing or deleted characters. In contrast, an extended gap prolongs an existing one, accommodating consecutive missing characters. This metric allows for smaller penalties for gap mismatches, resulting in more accurate measurements for truncated or shortened strings. Additionally, Jaro’s distance metric (Jaro, 1976) quantifies the similarity between two strings by identifying common characters that appear at the same positions in both strings and counting the number of transpositions. The Jaro metric is calculated using a formula that considers the number of shared characters, the lengths of the strings, and the number of transpositions. Monge et al. (Monge and Elkan, 1996) introduced a token-based similarity metric designed to detect duplicates in text fields using atomic strings. Atomic strings are identified by punctuation characters, acting as delimiters, and consist of alphanumeric characters as individual units within the text fields. Two atomic strings are considered duplicates if they are either identical or if one is a prefix of the other. This approach helps identify duplicates in text fields by considering matching atomic strings and provides a similarity score to assess the degree of duplication between fields.
@@
The Soundex algorithm, introduced by Russell et al. (Russell, 1922), is a phonetic coding scheme used to detect duplicates by comparing the phonetic similarity of surnames. The algorithm transforms surnames into codes based on rules of phonetic similarity. It preserves the initial letter of the surname as the prefix letter and assigns codes to each remaining letter according to specific phonetic groups. Vowels act as separators between consecutive consonants. Consecutive occurrences of the same code are consolidated, and if the resulting code has fewer than three characters, zeros are added as padding. By applying the Soundex algorithm, surnames are encoded into phonetic codes that capture their phonetic similarities, enabling the detection of similar-sounding surnames indicating potential duplicates.
Gupta et al. (Gupta et al., 2021) further enhance data readiness with their duplicate record detection metric in the toolkit. This metric aims to address the negative implications of data duplicates, such as resource wastage and data imbalance. Like other metrics in the toolkit, it provides a quality score ranging from 00 to $1$, with a score of $1$ indicating an absence of duplicates. The toolkit offers a remediation process that removes these duplicate records, ensuring clean data. Leveraging the “duplicated” function from the Pandas library (Wes McKinney, 2010), this metric provides a practical approach for identifying and resolving duplicate records.
@@
Summary: Duplicity refers to the presence of duplicate or redundant instances in a dataset, which can distort analysis and modeling. Bors et al. propose a uniqueness-based scoring system to flag duplicates. Various similarity metrics are introduced, including Levenshtein distance, Affine Gap Distance, Jaro’s distance, Monge et al.’s token-based algorithm, and the Soundex phonetic coding scheme. Gupta et al.’s toolkit provides a duplicate record detection metric that quantifies duplicate extent and offers a practical removal process.
4.1.5. Feature Relevancy

### v2

@@ 3. DRAI Metrics
This section provides an extensive summary of the existing metrics found in the literature that are used to measure data readiness for AI. We will explore these metrics for both structured and unstructured data, primarily focusing on structured data for which metrics have evolved more. Nevertheless, we describe many metrics related to assessing readiness of unstructured data, including textual, multimedia, image, speech, and video-related data. In Table 2, we provide a snapshot of all the dimensions and metrics discussed in this survey.
Table 2. Dimensions and Metrics for (Structured and Unstructured) DRAI
@@
Santos et al. (Santos et al., 2020), Pearson (Pearson, 2006)
@@
Templin (Templin, 1957), McCarthy et al. (McCarthy, 2005),
McCarthy et al.(McCarthy and Jarvis, 2010)
@@
Rousseeuw et al. (Rousseeuw and Hubert, 2018), Degirmenci et al. (Degirmenci and Karal, 2021)
@@
Mislabels Gupta et al’s (et al., 2021), Cohen’s Kappa (Cohen, 1960) Readability Score
@@
Monge et al. (Monge and Elkan, 1996), Russell et al. (Russell, 1922)
@@
Hall et al. (Hall and Smith, 1999)
@@
Gupta et al. (et al., 2021),Amazon SageMaker Developer Guide (Kemka, 2019)
@@
Leave-One-Out (Cook and Weisberg, 1982), Koh et al. (Koh and Liang, 2017),
Bachem et al. (Bachem et al., 2017), Ribeiro et al. (Ribeiro et al., 2016)
@@
Vatsalan et al. (Vatsalan et al., 2022), Duddu et al. (Duddu et al., 2022),
Carlini et al. (Carlini et al., 2022),Song et al. (Song and Mittal, 2021),
Bezzi (Bezzi, 2007), Longpr ’e et al. (Longpré et al., 2017),
Sevgi et al. (Arca and Hewett, 2020), Aindo AI (ain, nd)
@@
FAIR Compliance Wilkinson et al. (Wilkinson et al., 2018), Clarke et al. (Clarke et al., 2019)
@@ 3.1. Structured data
Structured data is organized with a consistent format and follows a specific structure or schema. Data stored in spreadsheets, relational database tables, self-describing file formats, etc., are common forms of structured data. This section will discuss metrics related to various dimensions, such as completeness, outliers, labels, etc., as shown on the left half of Figure 2 and their sources listed in the left half of Table 2.
@@ 3.1.1. Completeness
Example: In a dataset containing information about credit card customer demographics, this metric verifies if an attribute (e.g., income) is available for all customers. Completeness ensures that a dataset is reliable and suitable for analysis, as there is no loss of information due to missing data.
Metrics in Literature: Blake and Mangiameli (2011) propose a “completeness” metric to measure the presence of missing values in a dataset. This metric quantifies the proportion of null (missing) data records to the total number of data records.
Using this metric, researchers demonstrate how missing data can impact the results of classification tasks. For example, Jäger et al. (2021) demonstrates that handling missing values enhances predictive model performance. They observe up to $20\%$ improvement for classification tasks and $15\%$ improvement for regression tasks, emphasizing the importance of addressing missing data in optimizing downstream ML outcomes. In addressing missing data, Santos et al. (2020) use data imputation, which involves filling in or estimating missing values to maintain data integrity. In particular, the authors use the k-nearest neighbors (KNN) imputation technique in this process. They discuss the significance of choosing appropriate distance metrics, such as Heterogeneous Value Difference Metric (HVDM) and Heterogeneous Euclidean-Overlap Metri (HEOM), which effectively handle both nominal and continuous data while preserving data distribution during imputation.
Bors et al. (2018) propose a different approach to quantify missing values in a dataset by using indicators to distinguish missing from non-missing values. Their method offers a practical tool for data preparation, allowing easy identification of missing data in AI applications.
Another type of completeness targets missing data “disguised” with default values. Pearson (Pearson, 2006) discusses missing values are encoded or represented in ways that obscure their true nature, such as using “zero” values to indicate missing data. Such practices can severely distort analysis results. According to Vo et al. (2024) these issues affect not only the model’s predictions but also its explainability, potentially skewing feature importance calculations crucial for interpreting complex AI systems.
Impact on AI: Complete and accurate data enhances AI systems’ accuracy and reliability. Identifying explicitly missing data is easy, while disguised missing values pose a greater challenge as they appear valid but are placeholders or incorrect entries. Disguised values can lead to biased outcomes, reduced accuracy, and misinterpretation of AI models.
@@
Summary: Metrics proposed by various researchers quantify the impact of missing values and suggest remedies such as KNN imputation with suitable evaluation metrics. Additionally, using indicators for distinction and incorporating completeness metrics that consider data types and relationships further improve the handling of missing data.
@@ 3.1.2. Outliers
Outliers in a dataset refer to data points that significantly deviate from the typical or expected values within the dataset. They are points that are significantly distant from the majority of the data points and do not follow the general patterns present in the dataset.
@@
Example: Consider a dataset of housing prices based on factors such as size, number of bedrooms and bathrooms, location, schools, etc. In this dataset, an outlier could be a property with extremely high or low prices that do not align with the average price range of similar properties. This outlier might be an exceptional case, such as a luxury mansion in an otherwise average neighborhood, or it could be a data entry error.
@@
Metrics in Literature: In their research, Bors et al. (2018) discuss the concept of plausibility as a metric to identify outliers in datasets for AI applications, which can disrupt statistical analyses and modeling. Data analysts employ two main approaches to quantify the number of outliers: robust statistics and non-robust statistics. Robust statistics use the median and the robust interquartile range estimator, which are more resistant to outliers. Li et al. (2021) further explore the standard deviation and interquartile range-based outlier detection methods in their study. In contrast, non-robust statistics involve using the mean and standard deviation to identify entries that deviate significantly from the mean.
@@
In contrast, Breunig et al. (2000) introduce the Local Outlier Factor (LOF) as a metric for identifying outliers in a dataset. LOF quantifies the level of being an outlier for each data instance by considering the density of the dataset’s distribution. Outliers are expected to have lower local densities compared to their surrounding instances. The LOF algorithm computes the LOF value for a specific instance by comparing the density of that instance’s neighborhood with its neighboring instances. This neighborhood is a user-defined parameter (e.g., number of nearest neighbors). Higher LOF values indicate a higher probability of an instance being an outlier, which implies a notably lower density in its local neighborhood than in neighboring data points. Building on this foundation, Pokrajac et al. (2007) introduced the ILOF (Incremental Local Outlier Factor) method, which uses the LOF metric to determine if a new data point is an outlier. By analyzing the computed LOF value, the ILOF method assigns a score to the incoming sample, indicating whether it is classified as an outlier or not. This approach allows for real-time outlier detection and updates the scores of existing points to measure the impact of the new data point.
@@
Degirmenci and Karal (2021) introduce RiLOF, which addresses limitations in existing statistical outlier detection techniques by introducing the MoNNAD (Median of Nearest Neighborhood Absolute Deviation) metric. This metric is calculated as the median of the absolute variances among the LOF values of the $k$-nearest neighboring data points and the LOF value of a given sample. This score indicates how much the sample deviates from its local neighborhood. In the RiLOF method, the MoNNAD score is used to label and score query samples. Samples are categorized as outliers when their MoNNAD scores are equal to or greater than a specific limit. The RiLOF method assigns more importance to the query sample, resulting in clearer differentiation between inliers and outliers. The study demonstrates that the MoNNAD metric, incorporated in the RiLOF method, successfully detects outliers, including outlier clusters, that other techniques fail to recognize.
@@
The GESD (Generalized Extreme Studentized Deviation) technique, as introduced by Rosner (1983), and the MAD (Median Absolute Deviation) technique, proposed by Leys et al. (2013), are both outlier detection methods that share similarities in their underlying principles. Both approaches aim to identify outliers within datasets by using statistical measures to assess the deviation of data points from central tendencies. GESD identifies outliers by evaluating the maximum absolute difference between each sample and the dataset’s mean and normalizing it by the standard deviation. MAD identifies outliers by considering the median of absolute differences between data records and the dataset’s median. It incorporates a constant associated by assuming that data is normally distributed. Additionally, the Z-score method aligns with this principle by normalizing sample values using the mean and standard deviation or MAD. The robust Z-score version, introduced by Rousseeuw and Hubert (2018), substitutes the median and MAD for more robust measures, demonstrating the shared concept of using statistical measures to detect outliers in datasets.
Impact on AI: Outliers can significantly impact AI systems by skewing data distributions and introducing biases that lead to inaccurate models and unreliable predictions. They can distort statistical measures such as mean and variance, affecting algorithms like linear regression and clustering. They also can reduce the generalization and robustness of classification and neural network models. Outliers can also introduce noise, complicating the ML process and potentially leading to false positives or negatives (Infolabs, 2024). However, in specific contexts like fraud detection or rare disease diagnosis (e.g., Markham (2024)), outliers can be critical for identifying anomalies and should not be removed without a thorough analysis. Proper detection and management of outliers maintain the integrity and effectiveness of AI models.
@@
Summary: Outliers data points deviate significantly from most of the dataset and do not follow the general patterns. Different metrics and techniques proposed to identify and measure outliers in datasets include measures based on column heterogeneity, statistical measures like median, standard deviation, mean, and interquartile range, and techniques such as Local Outlier Factor (LOF), Generalized Extreme Studentized Deviation (GESD), Z-score.
@@ 3.1.3. Mislabeled Data
Mislabeled data in the context of preparing a dataset for AI refers to instances or data points with inaccurate labels. It represents a form of labeling error or inconsistency within the dataset, where the assigned labels do not align with the true or expected labels.
@@
Metrics in Literature: Gupta et al.’s (et al., 2021) Data Quality Toolkit (DQT) introduces a label purity metric to measure the impact of adding random noise on the performance of a classifier trained on the dataset. In the example provided in the study, $10\%$ random noise is introduced to $41$ datasets from UCI (Dua and Graff (2017)) and Kaggle ((Kag, nd)) repositories, and the performance of an AutoAI classifier (Liu et al. (2020)) is measured using 3-fold cross-validation. The results show a drop in classifier performance after inducing noise, with varying degrees of decrease observed across the datasets.
In evaluating the accuracy of labels assigned by multiple annotators, Cohen’s Kappa (Cohen, 1960) is widely employed for assessing inter-rater reliability, especially with categorical or binary labels. Cohen’s Kappa calculates the agreement beyond chance, ranging from $-1$ to $1$, where values near $1$ indicate substantial agreement. Lavitas et al. (2021) contribute a credibility metric, assessing the likelihood of correct annotations based on multiple reviewers’ agreement. The metric ranges from $(N/2+1)/N$ to $1$, reflecting high credibility with close agreement and lower credibility with less agreement among reviewers, offering insights into the reliability of the annotation process involving multiple reviewers
Impact on AI: Incorrect labels in training data lead to poor model performance because the AI learns from erroneous examples, resulting in skewed predictions and decisions (Cleanlab, 2024). This issue can cause substantial financial costs due to the need to retrain models and correct errors. Additionally, mislabeled data can undermine trust in AI systems, as stakeholders lose confidence in their outputs. Ethical implications are also significant, as mislabeled data can introduce or amplify biases that lead to discriminatory outcomes, such as flawed medical diagnoses or biased hiring algorithms. Addressing mislabeled data is crucial for maintaining the integrity and effectiveness of AI models.
@@ 3.1.4. Duplicate Values
This refers to the presence of duplicate or redundant records within a dataset. Duplicates appear when the same or similar data entries appear multiple times, potentially skewing the ML process.
Example: Consider a dataset of customer transactions where each entry represents a purchase made by a customer. Multiple entries of the same transaction refer to duplication, which may impact the analysis results.
@@
Metrics in Literature: Bors et al. (2018) propose a mechanism to identify duplicate entries in a dataset using a scoring system based on uniqueness. A user selects one or more columns in the dataset that are intended to have distinctive combinations of values. The system assigns a score of $1$ (true) to values with a unique combination in the selected columns and a 00 (false) score to values found multiple times. By incorporating this score, the system effectively flags duplicate entries in the dataset. Elmagarmid et al. (2007) conducted a comprehensive survey exploring various similarity metrics for duplicate detection, addressing challenges in managing typographical variations in string data. One of the highlighted character-based similarity metrics in the survey is the Levenshtein distance metric (Levenshtein, 1965). This metric measures the number of operations needed to transform one string into another through edit operations (insertion, deletion, and character replacement). Waterman et al. (1976) introduced the Affine Gap Distance metric to overcome the limitations of the standard edit distance metric in matching shortened or truncated strings. It introduces two edit operations: open and extend the gap. An open gap in sequence alignment indicates the start of missing or deleted characters in the sequence. In contrast, an extended gap accommodates consecutive missing characters by extending an existing one. This metric allows for smaller penalties for gap mismatches, resulting in more accurate measurements for truncated or shortened strings. Additionally, Jaro’s distance metric (Jaro, 1976) quantifies the similarity between two strings by identifying common characters that appear at the same positions in both strings and adding up the number of transpositions. The Jaro metric considers the number of shared characters, the lengths of the strings, and the number of transpositions. Monge and Elkan (1996) introduced a token-based similarity metric designed to detect duplicates in text fields using atomic strings. Atomic strings are identified by punctuation characters, acting as delimiters, and consist of alphanumeric characters as individual units within the text fields. Two atomic strings are considered duplicates if they are either identical or if one is a prefix of the other. This approach helps identify duplicates in text fields by considering matching atomic strings and provides a similarity score to assess the degree of duplication between fields.
@@
The Soundex algorithm, introduced by Russell (1922), is a phonetic coding scheme used to detect duplicates by comparing the phonetic similarity of character strings, such as names. The algorithm transforms names into codes based on rules of phonetic similarity. It preserves the initial letter of the name as the prefix letter and assigns codes to each remaining letter according to specific phonetic groups. Vowels act as separators between consecutive consonants. Consecutive occurrences of the same code are merged, and if the resulting code has fewer than three characters, zeros are added as padding. By applying the Soundex algorithm, names are encoded into phonetic codes that capture their phonetic similarities. It enables the detection of similar-sounding names indicating potential duplicates.
Impact on AI: When training on duplicated data, models may overfit by learning redundant patterns that do not generalize well to unseen data. This over-representation can lead to skewed predictions and unreliable outcomes. Duplicates also increase the dataset size, storage costs, and computational resources required for training. Additionally, they can degrade data quality, making it harder to derive accurate insights and leading to inefficiencies in data processing.
@@
Summary: Duplicates refer to the presence of duplicate or redundant instances in a dataset, which can distort analysis and modeling. Various similarity metrics, based on Levenshtein distance, Affine Gap Distance, Jaro’s distance, Monge et al.’s token-based algorithm, and the Soundex phonetic coding scheme, are available to measure duplicates.
3.1.5. Feature Relevance